OpenStudy (lilmike234):

which of the following graphs matches the circle defined by this equation (x-3)^2+(y+1)^2=25

OpenStudy (usukidoll):

The center-radius form of the circle equation is in the format (x – h)^2 + (y – k)^2 = r^2, with the center being at the point (h, k) and the radius being "r".

OpenStudy (usukidoll):

we just need tiny bit of information to graph this...

OpenStudy (usukidoll):

\[ (x-3)^2+(y+1)^2=25 \] what is the square root of 25?

OpenStudy (lilmike234):

5

OpenStudy (lilmike234):

Here are the answer choices

OpenStudy (usukidoll):

ok, and what will our h and k be? hint: opposite signs for h,k . it's similar to solving for x and y

OpenStudy (lilmike234):

(-2,8) ?

OpenStudy (usukidoll):

where did that come from? O_O

OpenStudy (triciaal):

@Lilmike234 focus on the process the answer will come then you match with what you get

OpenStudy (lilmike234):

Im having a hard time finding the answer but I think it was answer choice B

OpenStudy (triciaal):

just like the other problem finding the value of a

OpenStudy (usukidoll):

no... that's not it. It's not B.

OpenStudy (usukidoll):

what is the center for the equation (x-3)^2+(y+1)^2=25 ?

OpenStudy (lilmike234):

(-3,1)=5

OpenStudy (usukidoll):

opposite signs for the center

OpenStudy (usukidoll):

x-3= 0 y+1 = 0 solve for x and y

OpenStudy (lilmike234):

x=3 y=-1

OpenStudy (usukidoll):

ok now center is (3,-1) and radius is 5

OpenStudy (lilmike234):

So my answer would be C right ?

OpenStudy (usukidoll):

 that's where our center point is